OpenStudy (anonymous):

what is the equation of the line passing through (-3,-9) and parallel to the line y=x+6 in standard form?

OpenStudy (anonymous):

so if y=mx+b... then this slope is 1 and the y int is 6. since the lines are parallel, the slopes are identical.. thus y-y1 = m(x-x1) to find your equation. It is really simple if you just try to do it. you know that -3 is the x1 coordinate and -9 is the y1 coordinate that are given.. thus you plug in the numbers.. y-(-9) = 1(x-(-3)) - solve for y

OpenStudy (anonymous):

I got x-y=6 final answer

OpenStudy (anonymous):

i got y = x - 6